OpenStudy (anonymous):

Determine the mass of 2.5 cups of water if the density of water is 1.00g/cm^3 and 1 cup=240mL.

OpenStudy (abmon98):

1 cup=240cm^3 2.5 cups= x cm^3 Cross multiplication to solve for x. x=2.5*240/1=600 cm^3 Density=mass(g)/Volume(cm^3) 1.00=mass(g)/600 600*1=Mass(g)

OpenStudy (anonymous):

Thank you so much! :)

OpenStudy (abmon98):

your most welcome :D